Calculate λmax for blackbody radiation for the following.

(a) liquid helium (3.8 K)

(b) room temperature (293 K)

(c) a steel furnace (1200 K)

(d) a blue star (9200 K)
Answer #1

λmax*T= 2.89*10^-3 mK

λmax=2.89*10^-3 mK/3.8

=0.760*10^-3 m---a

b)λmax*T= 2.89*10^-3 mK

λmax= 2.89*10^-3 mK/293

=0.00986*10^-3 m--b

c)

λmax*T= 2.89*10^-3 mK

λmax= 2.89*10^-3 mK/1200

=0.0024*10^-3 m--c

d)

λmax*T= 2.89*10^-3 mK

λmax= 2.89*10^-3 mK/9200

=0.00031*10^-3 m